OpenStudy (anonymous):

inverse variation help 1.suppose that y varies inversely with x, and y=6 when x=8. what is an equation for the inverse variation? 2.suppose that y varies inversely with x, and y=0.2 when x=2. what is an equation for the inverse variation?

OpenStudy (anonymous):

1. options y=x/48 y=2x x=y/2 y=48/x 2.options y=1.8x y=0.4/x x=y/1.8 y=x/0.4

OpenStudy (anonymous):

do you know what "y varies inversely with x" means?

OpenStudy (anonymous):

i know what inverse is but no I'm not sure

OpenStudy (anonymous):

then i will tell you it means there is some number, usually denoted by \(k\) but it is just some number, where \[\huge y=\frac{k}{x}\] and your job is to find that \(k\)

OpenStudy (anonymous):

take the expression \(\huge y=\frac{k}{x}\) where you see a \(y\) put \(6\) where you see an \(x\) put \(8\) you will have a simple equation to solve for \(k\)

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

i am holding...

OpenStudy (anonymous):

i got 48

OpenStudy (anonymous):

so 48/x is the correct answer?

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

yeah \[y=\frac{48}{x}\]

OpenStudy (anonymous):

and 2 is .4/x :)

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

okay so 2 more questions or 1 really how do i find out if a table represents a direct variation or inverse

OpenStudy (anonymous):

whats the whole question

OpenStudy (anonymous):

NICE WORK

OpenStudy (anonymous):

and what that variation is

OpenStudy (anonymous):

it is nice when someone actually tries something rather than repeatedly asking the same questions

OpenStudy (anonymous):

thank you for helping me(: okay this is a table x:2,4,6,8 7:14,28,42,56

OpenStudy (anonymous):

you have to look at two points on the table or graph

OpenStudy (anonymous):

that is your table, it is pretty clear that as x gets bigger, so does y, so it COULD be direct variation

OpenStudy (anonymous):

you can check it by checking to see if the first number, the x, is multiplied by the same number each time to get the y

OpenStudy (anonymous):

everything is times 7 isnt that the direct varition

OpenStudy (anonymous):

i think its direct and y=7x?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

haha you typed that before me

OpenStudy (anonymous):

you can usually do it in your head, as in this case if it is not obvious check \(\frac{y}{x}\) and see if you keep getting the same number

OpenStudy (anonymous):

if, as x gets bigger, y gets smaller it MAY be inverse variation or whatever they call it

OpenStudy (anonymous):

I'm looking at this one and its not as easy x:2,4,8,12 y:4,2,1,2/3

OpenStudy (anonymous):

it is always possible the answer is NO

OpenStudy (anonymous):

so this is inverse because the y gets smaller

OpenStudy (anonymous):

here as x gets bigger y gets smaller so check to see if the product of the two numbers is always the same

OpenStudy (anonymous):

yea right and what do you get when you multiply each pair of numbers?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

you should see that in fact this is even easier than checking direct variation right easy yes?

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

so you know what it looks like as an equation right?

OpenStudy (anonymous):

yes thank you so much

OpenStudy (anonymous):

yw